# Sequential Randomization Tests Using e-values: Applications for trial monitoring

### v5

@@ Abstract
Sequential monitoring of randomized trials traditionally relies on parametric assumptions or asymptotic approximations. We discuss a nonparametric sequential test and its application to continuous and time-to-event endpoints that derives validity solely from the randomization mechanism. Using a betting framework, these tests constructs a test martingale by sequentially wagering on treatment assignments given observed outcomes. Under the null hypothesis of no treatment effect, the expected wealth cannot grow, guaranteeing anytime-valid Type I error control regardless of stopping rule. We prove validity and present simulation studies demonstrating calibration and power. These methods provide a conservative, assumption-free complement to model-based sequential analyses.
@@ 1 Introduction
E-values and e-processes offer an alternative framework (shafer2021; vovk2021; ramdas2025). An e-value is a measure of evidence against a null hypothesis with a specific property: its expected value under the null is at most 1. This simple constraint yields anytime-valid inference: the Type I error guarantee holds at any stopping time, regardless of the stopping rule.
pmlr-v177-duan22a introduced interactive rank testing by betting (i-bet), which tests treatment effects by wagering on treatment assignments given observed outcomes. The intuition is discussed by ramdas2021game. Under the null hypothesis, randomization ensures that assignments are independent of outcomes, so no betting strategy can systematically accumulate wealth.
We propose e-RT (e-value Randomized Trial), a method for prospective sequential monitoring of randomized trials. Like i-bet (pmlr-v177-duan22a), e-RT uses betting martingales for inference, but differs in key respects: e-RT monitors sequentially as patients enroll rather than analyzing completed trial data; it requires no covariates or working models; and it uses an adaptive betting fraction derived from running outcome estimates rather than fixed wagers. This yields a method with minimal assumptions suitable for real-time trial monitoring.
@@ 2.2 Wealth Process
Following pmlr-v177-duan22a, we construct a wealth process by wagering on treatment assignments. After observing outcome $Y_{i}$ but before learning treatment assignment $T_{i}$, we choose $\lambda_{i}\in[0,1]$: the fraction wagered on intervention.
@@ 2.5 Validity
By Ville’s inequality (ville1939), for any nonnegative martingale starting at 1:
@@ 4.2 Betting strategy for continuous outcomes
$\lambda_{i}=0.5+c_{i}c_{\max}g_{i}.$ (14)
By construction, $\lambda_{i}\in(0,1)$ and is predictable: it depends only on past outcomes and the new $Y_{i}$, not on $T_{i}$.
@@
  • If $Y_{i}$ is close to the historical median, $g_{i}\approx 0$ and $\lambda_{i}\approx 0.5$: we essentially do not bet.
  • If $Y_{i}$ is unusually low (for example, very high ventilator-free days in a trial where we expect intervention to improve outcomes), then $g_{i}<0$ or $g_{i}>0$ depending on how we code the outcome and the expected direction of benefit. The sign of $g_{i}$ determines which arm we lean towards.
@@ 4.4 Worked intuition
Imagine a trial where the outcome is ventilator-free days, and higher is better. Suppose that after 100 patients, the median and MAD of $Y$ are roughly stable. Patient 101 has an unusually high number of ventilator-free days compared with this distribution. The standardized residual $r_{101}$ is positive and large, so $g_{101}\approx 0.8$ and, after burn-in, $c_{101}\approx 1$. With $c_{\max}=0.6$, we might get $\lambda_{101}\approx 0.5+0.6\times 0.8\approx 0.98$: we strongly bet that this patient was in the intervention arm. If they indeed were, wealth increases by roughly a factor of $0.98/0.5\approx 2$ for this one patient. If not, wealth shrinks by $(1-0.98)/0.5\approx 0.04$.
@@ 4.6 Simulation overview
The test was run sequentially with a burn-in period and ramp (for example, burn-in = 20 patients, ramp = 50 patients, $c_{\max}=0.6$). Under the null (no mean difference between arms), Type I error was close to or below the nominal level. Under the alternative, the implementation for continuous endpoints rejected the null with moderate power and typically crossed the threshold at an intermediate sample size (Table 2). Again, we believe that this may not be a replacement for the fixed-sample $t$-test but a conservative, anytime-valid monitoring tool that can trigger early stopping when effects are larger or clearer than anticipated.
@@ 5 Time-to-Event Outcomes
pmlr-v146-grunwald21a developed a safe logrank test using e-values under a proportional hazards model with a prior on the hazard ratio. We attempted to construct a nonparametric approach where validity derives solely from randomization, not from a correctly specified hazard model. We call this e-survival.
@@ 5.2 Betting Strategy
Note that unlike the binary and continuous approaches, this betting strategy uses only the sign of the cumulative score $Z_{j-1}$, not its magnitude. Once evidence favors one direction, the bet size is fixed at $\lambda_{\max}$ regardless of how strong the accumulated evidence is. This mirrors the Kelly criterion in betting theory: the optimal wager size depends on the expected edge, and $\lambda_{\max}=0.25$ is calibrated for moderate effects (HR $\approx 0.80$(kelly1956). The connection between the wager, $\lambda$, and Kelly’s ideas on fraction of betting needs to be further explored. In brief, it makes sense that wager should be higher when prospects of winning are more favorable. For time-to-event, however, adapting response to events may take a long time, and a fixed wager may be preferable.

### v6

@@ Abstract
Sequential monitoring of randomized trials traditionally relies on parametric assumptions or asymptotic approximations. We discuss a family of nonparametric sequential tests—collectively called e-RT—for binary, deaths-only, continuous, time-to-event, and multi-state endpoints. All variants derive validity solely from the randomization mechanism. Using a betting framework, each test constructs a test martingale by sequentially wagering on treatment assignments given observed outcomes. Under the null hypothesis of no treatment effect, the expected wealth cannot grow, guaranteeing anytime-valid Type I error control regardless of stopping rule. We prove validity for each variant, present simulation studies demonstrating calibration and power, and discuss the principled asymmetry in betting strategies across outcome types. These methods provide a conservative, assumption-free complement to model-based sequential analyses.
@@ 1 Introduction
E-values and e-processes offer an alternative framework (Shafer, 2021; Vovk and Wang, 2021; Ramdas and Wang, 2025). An e-value is a measure of evidence against a null hypothesis with a specific property: its expected value under the null is at most 1. This simple constraint yields anytime-valid inference: the Type I error guarantee holds at any stopping time, regardless of the stopping rule.
Duan et al. (2022) introduced interactive rank testing by betting (i-bet), which tests treatment effects by wagering on treatment assignments given observed outcomes. The intuition is discussed by Ramdas (2021). Under the null hypothesis, randomization ensures that assignments are independent of outcomes, so no betting strategy can systematically accumulate wealth.
We propose e-RT (e-value Randomized Trial), a family of methods for prospective sequential monitoring of randomized trials. Like i-bet (Duan et al., 2022), e-RT uses betting martingales for inference, but differs in key respects: e-RT monitors sequentially as patients enroll rather than analyzing completed trial data; it requires no covariates or working models; and it uses adaptive betting fractions derived from running outcome estimates rather than fixed wagers. This yields methods with minimal assumptions suitable for real-time trial monitoring.
We present five variants: the binary e-RT for event/no-event outcomes; e-RTd for deaths-only monitoring (requiring no survivor tracking); e-RTc for continuous endpoints; e-RTs for time-to-event data; and e-RTms for multi-state trajectory data. All share the same validity proof—the expected wealth multiplier is exactly 1 under the null—but differ in how they translate outcome data into wagers. We also discuss the principled asymmetry in betting strategy design: why sparse-update methods (survival, deaths-only) can afford aggressive fixed wagers while dense-update methods (binary, continuous) require conservative adaptive strategies.
@@ 2.2 Wealth Process
Following Duan et al. (2022), we construct a wealth process by wagering on treatment assignments. After observing outcome $Y_{i}$ but before learning treatment assignment $T_{i}$, we choose $\lambda_{i}\in[0,1]$: the fraction wagered on intervention.
@@ 2.5 Validity
By Ville’s inequality (Ville, 1939), for any nonnegative martingale starting at 1:
@@ 3.3 Trajectory Examples
4 Deaths-Only Monitoring (e-RTd)
4.1 Motivation
The binary e-RT requires knowing both the treatment arm and the outcome for every enrolled patient. In practice, this means a coordinator must ascertain whether each patient experienced the event or not—requiring follow-up, data entry, and outcome adjudication for all patients, including the majority who survive.
In some settings, especially those embedded in electronic medical records (EMR) or when studying mortality in acutely ill patients, deaths are reliably captured but survivors may not be systematically followed. A death is an unmissable event; a survivor at day 28 requires active confirmation. This motivates a simpler variant: e-RTd monitors only the stream of deaths, ignoring survivors entirely.
The trade-off is explicit: by discarding information about survivors, we lose statistical efficiency (approximately $2.5\times$ sample size inflation compared with frequentist methods). What we gain is operational simplicity—the only data required are the arm labels on deaths, in the order they occur. No denominators, no follow-up windows, no survivor tracking.
4.2 Null Hypothesis
Consider a trial with 1:1 randomization. Under the null hypothesis of no treatment effect on mortality, both arms have the same mortality rate. If both arms have equal mortality and equal enrollment, then each death is equally likely to come from either arm:
$H_{0}:P(\text{death from treatment}\mid\text{a death occurred})=0.5$ (10)
In plain language: if the treatment does not affect mortality, then looking at the arm label on a death certificate is like flipping a fair coin. Treatment deaths and control deaths should arrive in roughly equal numbers.
This reduces trial monitoring to a sequential test of a Bernoulli coin: is the fraction of treatment deaths equal to $0.5$? We call this the “death coin.”
4.3 Algorithm
The algorithm maintains two counters: $d_{\text{trt}}$ and $d_{\text{ctrl}}$, both initialized to zero. For each death $i=1,2,\ldots$:
Step 1: Estimate the coin bias from past data. Using all deaths before the current one, compute the plug-in estimator:
$\hat{p}_{i-1}=\frac{d_{\text{trt}}}{d_{\text{trt}}+d_{\text{ctrl}}}$ (11)
If no deaths have been observed yet, set $\hat{p}=0.5$. In words: $\hat{p}$ is the running fraction of deaths that came from the treatment arm. Under the null, this should hover around $0.5$. Under the alternative (if treatment helps), this should drift below $0.5$.
Step 2: Compute the wager. The betting fraction is:
$\lambda_{i}=0.5+c_{i}\cdot(\hat{p}_{i-1}-0.5)$ (12)
clamped to $[0.001,0.999]$, where $c_{i}$ is the same ramp function as before:
$c_{i}=\min\left(1,\max\left(0,\frac{i-n_{0}}{n_{r}}\right)\right)$ (13)
with burn-in $n_{0}=30$ deaths and ramp $n_{r}=50$ deaths. During the burn-in, $c_{i}=0$ and $\lambda_{i}=0.5$ (a neutral bet—no information yet). After the ramp completes, $c_{i}=1$ and $\lambda_{i}=\hat{p}_{i-1}$: the wager fully reflects the observed death proportion.
In plain language: we bet proportionally to what we have learned so far. If 40% of past deaths were from treatment ($\hat{p}=0.40$), we set $\lambda=0.40$. This means we place more of our wager on the “control death” side, because past data suggest control deaths are more common.
At full ramp ($c_{i}=1$), this is the full Kelly bet—the wager that maximizes the expected logarithmic growth rate of wealth. This contrasts with the half-Kelly strategy used in the binary e-RT. Full Kelly is viable here because events are sparse: a trial with 2,000 patients and 15% mortality produces approximately 300 deaths. Over-betting compounds across only a few hundred multiplications, causing gradual wealth erosion rather than the catastrophic collapse seen with dense (every-patient) updates.
Step 3: Update wealth.
[…]
starting from $W_{0}=1$.
In plain language: we place $\lambda_{i}$ on “treatment death” and $1-\lambda_{i}$ on “control death.” The null hypothesis pays $2:1$ on the correct side (a fair payout for a $0.5$-probability event). If we guessed correctly, wealth grows; if we guessed wrong, wealth shrinks.
Step 4: Update counters after betting. Increment $d_{\text{trt}}$ or $d_{\text{ctrl}}$ depending on which arm the death came from. This ordering—bet first, then update—ensures that the estimate $\hat{p}_{i-1}$ uses only past information, maintaining the martingale property.
4.4 Worked Example
Consider a trial comparing a new treatment against standard of care in the ICU. Suppose 80 deaths have been observed so far. Burn-in (30) and ramp (50) are complete, so $c_{i}=1$. Current counts: $d_{\text{trt}}=33$, $d_{\text{ctrl}}=47$.
Death 81 arrives. Let us walk through the update.
Step 1: $\hat{p}_{80}=33/(33+47)=33/80=0.4125$. Interpretation: so far, 41.25% of deaths came from the treatment arm. This is below 50%, suggesting treatment may be protective.
Step 2: $\lambda_{81}=0.5+1.0\times(0.4125-0.5)=0.4125$. We place 41.25% of our bet on “treatment death” and 58.75% on “control death.” We are leaning toward control because past evidence suggests treatment deaths are less frequent.
Step 3: The arm is revealed: it is a control death.
$\text{Multiplier}=\frac{1-0.4125}{0.5}=\frac{0.5875}{0.5}=1.175$ (15)
Wealth grows by 17.5%. Our bet was correct—we leaned toward control, and indeed it was a control death.
Step 4: Update $d_{\text{ctrl}}=48$. Now $\hat{p}_{81}=33/81=0.407$.
Had the death been from the treatment arm instead:
$\text{Multiplier}=\frac{0.4125}{0.5}=0.825$ (16)
Wealth would have shrunk by 17.5%. Our lean toward control would have been wrong.
Under the null, treatment and control deaths arrive with equal probability, so wins and losses balance on average. Under the alternative ($\hat{p}<0.5$), control deaths are genuinely more frequent, and wealth grows systematically.
4.5 Validity
Theorem 2.
[…]
Proof.
Condition on all past information and the current wager $\lambda_{i}$. Under the null, the probability that death $i$ is from the treatment arm is exactly $0.5$, independently of all past data. The expected wealth multiplier is:
$\mathbb{E}\left[\frac{W_{i}}{W_{i-1}}\,\middle|\,\lambda_{i}\right]$ $=0.5\times\frac{\lambda_{i}}{0.5}+0.5\times\frac{1-\lambda_{i}}{0.5}$ (17)
$=\lambda_{i}+(1-\lambda_{i})=1$ (18)
This is the same identity as in the binary case. Since the expected multiplier is exactly 1, wealth cannot systematically grow under the null, regardless of how $\lambda_{i}$ was chosen (as long as it depends only on past deaths). ∎
By Ville’s inequality, $\Pr_{H_{0}}(\sup_{i\geq 1}W_{i}\geq 1/\alpha)\leq\alpha$. Rejecting when wealth crosses $1/\alpha$ controls Type I error at any stopping time.
4.6 Bidirectionality
The method automatically detects both benefit and harm without pre-specifying a direction:
  • If $\hat{p}<0.5$ (fewer treatment deaths than expected): the method bets on control deaths being more common, and wealth grows when treatment is indeed protective.
  • If $\hat{p}>0.5$ (more treatment deaths than expected): the method bets on treatment deaths being more common, and wealth grows when treatment is harmful.
No investigator input about the expected direction is needed. The adaptive wager discovers the direction from the data.
4.7 Signal Concentration
A key property of e-RTd is that it can outperform the full-sample binary e-RT when baseline mortality is low. The intuition is that deaths concentrate the treatment signal.
Consider a trial where the control mortality is 25% and treatment mortality is 20%, yielding a 5 percentage-point absolute risk reduction (ARR). In the full-sample binary e-RT, the signal per patient is diluted: most patients survive, and only the 20–25% who die carry information about differential mortality. The observed risk difference across all patients is 5 percentage points.
In e-RTd, only deaths are observed. The death-coin probability is:
$p_{\text{alt}}=\frac{p_{\text{trt}}}{p_{\text{trt}}+p_{\text{ctrl}}}=\frac{0.20}{0.20+0.25}=0.444$ (19)
This is an 11.2-point tilt from 0.5—more than double the 5-point ARR. The signal is concentrated because deaths filter out the uninformative survivors.
This advantage diminishes as baseline mortality increases. At higher event rates, the binary method sees more informative events per patient, and the death-coin tilt shrinks because both numerator and denominator grow:
Table 2: Signal Concentration: Death-Coin Tilt vs. ARR for a 5pp Risk Reduction
Baseline Mortality Treatment Mortality Death Coin $p_{\text{alt}}$ Tilt from 0.5 Tilt / ARR
[…]
The crossover with the binary e-RT occurs at approximately 25% baseline event rate (Table 4). Below that, the concentrated death-coin signal more than compensates for the smaller number of observations; above it, the binary method’s access to all patients provides the advantage.
4.8 Simulation Studies
We evaluated e-RTd operating characteristics using the same simulation framework as for binary e-RT. For each scenario, we calculated the frequentist sample size (two-proportion $z$-test) and applied a $2.5\times$ inflation factor to account for the reduced efficiency of deaths-only monitoring. We used burn-in =30=30 deaths, ramp =50=50 deaths, threshold $1/\alpha=20$, and 2,000 simulations per scenario.
Table 3: Operating Characteristics for e-RTd (Deaths-Only)
Baseline ARR Death Coin N (freq) N (e-RTd) Type I Power Median Crossing
15% 5pp 0.400 1,372 3,430 2.20% 90.3% 179 deaths (42%)
15% 10pp 0.250 282 705 0.35% 59.6% 63 deaths (89%)
25% 5pp 0.444 2,188 5,470 2.90% 86.0% 490 deaths (40%)
25% 10pp 0.375 500 1,250 2.15% 87.1% 123 deaths (49%)
35% 5pp 0.462 2,754 6,885 3.05% 76.1% 1,036 deaths (46%)
35% 10pp 0.417 658 1,645 3.15% 81.1% 232 deaths (47%)
Type I error was well controlled across all scenarios (0.35–3.15%), consistently below the nominal $\alpha=0.05$. Power ranged from 59.6% to 90.3%, with the strongest performance at low baseline mortality rates where the signal concentration effect is most pronounced: at 15% baseline mortality with a 5pp ARR, the death-coin tilt of 10 points from 0.5 yielded 90.3% power. At higher baseline rates (35%), the death-coin tilt shrinks and power decreases, consistent with the signal concentration analysis above.
When the threshold was crossed, it occurred at approximately 40–50% of total expected deaths for most scenarios. The exception is 15% baseline with 10pp ARR, where the very strong death-coin tilt (0.250, a 25-point deviation from 0.5) drives rapid crossing at the cost of requiring very few deaths overall (median 63), so the 89% figure reflects that the total expected death count (71) is small and crossing happens close to the end.
4.9 Trajectory Examples
Figure 3 shows example wealth trajectories for e-RTd. Under the null (left), the death coin is fair and wealth fluctuates around 1. Under the alternative (right), treatment deaths arrive less frequently than control deaths, the adaptive wager learns this imbalance, and wealth grows.
Figure 3: Trajectories of the e-RTd process (25% baseline mortality, 5pp ARR, 500 deaths). Left: under the null hypothesis (death coin =0.50=0.50), wealth fluctuates randomly. Right: under the alternative hypothesis, wealth grows as the adaptive wager learns the death-coin imbalance. Dashed red line: rejection threshold ($1/\alpha=20$).
4.10 Head-to-Head Comparison with Binary e-RT
To quantify the signal concentration crossover, we ran both e-RTd and binary e-RT on the same simulated trials across a range of baseline mortality rates with a fixed 5pp ARR. For each baseline rate, we computed the frequentist sample size (two-proportion $z$-test, 80% power), enrolled that many patients, and analyzed the data with both methods: binary e-RT processed all patients; e-RTd processed only the death stream. We used 2,000 simulations per scenario.
Table 4: Head-to-Head Comparison: e-RTd vs. Binary e-RT (5pp ARR, same trial data)
Baseline Death Coin NN Deaths Binary Power e-RTd Power $\Delta$ Winner
10% 0.333 870 66 12.2% 8.4% 3.7-3.7pp Binary
15% 0.400 1,372 172 26.7% 44.6% +17.9+17.9pp e-RTd
20% 0.429 1,812 318 32.4% 44.0% +11.6+11.6pp e-RTd
[…]
30% 0.455 2,502 689 42.3% 38.4% 3.9-3.9pp Binary
35% 0.462 2,754 896 45.3% 36.1% 9.2-9.2pp Binary
40% 0.467 2,942 1,104 48.8% 33.1% 15.6-15.6pp Binary
The crossover occurs at exactly 25% baseline mortality, confirming the analytical prediction from Table 2. At 15–20% baseline, e-RTd outperforms binary e-RT by 12–18 percentage points despite seeing fewer observations, because the death-coin tilt (10–7 points from 0.5) more than compensates for the smaller sample. Above 30%, binary e-RT’s access to all patients provides an increasingly large advantage.
The 10% exception.
At 10% baseline, the frequentist $N=870$ produces only $\sim$66 expected deaths—fewer than the burn-in (30) plus ramp (50) = 80 deaths required for e-RTd to reach full betting strength. The e-RTd process literally never completes its learning phase, so it cannot leverage the strong 16.7-point death-coin tilt. This highlights a practical constraint: e-RTd requires sufficient deaths (>80>80) to be effective.
The 10pp ARR case.
With a larger effect (10pp ARR), the frequentist $N$ shrinks dramatically (282 at 15% baseline), producing even fewer deaths. Binary e-RT dominates across all tested baseline rates because the larger per-patient effect can be detected with fewer observations, while e-RTd cannot accumulate enough deaths during its burn-in/ramp phase.
Practical guidance.
Use e-RTd when (i) baseline mortality is 15–25%, (ii) the expected ARR is modest ($\leq$5pp), and (iii) outcome ascertainment beyond death is impractical. Above 25% baseline or with larger expected effects, binary e-RT is more powerful.
@@ 5.2 Betting strategy for continuous outcomes
Additionally, we need a direction estimate: which arm has better outcomes? Using all previous data, we compute the running Cohen’s $d$:
$\hat{d}_{i-1}=\frac{\bar{Y}_{\text{trt}}-\bar{Y}_{\text{ctrl}}}{s_{\text{pooled}}}$ (24)
clamped to [1,1][-1,1], where $\bar{Y}_{\text{trt}}$ and $\bar{Y}_{\text{ctrl}}$ are the arm-specific means and $s_{\text{pooled}}$ is the pooled standard deviation. This is the “doubly adaptive” structure: $g_{i}$ captures how informative the current observation is, while $\hat{d}_{i-1}$ captures the direction and magnitude of the treatment effect estimated from all past data.
@@
$\lambda_{i}=0.5+c_{i}\cdot c_{\max}\cdot g_{i}\cdot\hat{d}_{i-1}.$ (25)
By construction, $\lambda_{i}\in(0,1)$ and is predictable: it depends only on past outcomes and the new $Y_{i}$, not on $T_{i}$. In words: if the current observation $Y_{i}$ is extreme in the direction that past data associate with the treatment arm, $\lambda_{i}$ deviates substantially from 0.5, placing a confident bet on treatment. If $Y_{i}$ is typical, or if past data show no treatment effect ($\hat{d}\approx 0$), $\lambda_{i}$ stays near 0.5 and we barely bet.
@@
  • If $Y_{i}$ is close to the historical median, $g_{i}\approx 0$ and $\lambda_{i}\approx 0.5$: we essentially do not bet, regardless of how strong the estimated treatment effect is.
[…]
  • If $Y_{i}$ is extreme but $\hat{d}\approx 0$ (no estimated effect yet), $\lambda_{i}$ stays near 0.5—an unusual observation is not informative if we do not yet know which direction to bet.
@@ 5.4 Worked intuition
Imagine a trial where the outcome is ventilator-free days, and higher is better. Suppose that after 100 patients, the median and MAD of $Y$ are roughly stable, and the running Cohen’s $d$ estimate is approximately 1.0 (a large effect). Patient 101 has an unusually high number of ventilator-free days compared with this distribution. The standardized residual $r_{101}$ is positive and large, so $g_{101}\approx 0.8$ and, after burn-in, $c_{101}\approx 1$. With $c_{\max}=0.6$ and $\hat{d}\approx 1.0$, we get $\lambda_{101}\approx 0.5+0.6\times 0.8\times 1.0\approx 0.98$: we strongly bet that this patient was in the intervention arm. If they indeed were, wealth increases by roughly a factor of $0.98/0.5\approx 2$ for this one patient. If not, wealth shrinks by $(1-0.98)/0.5\approx 0.04$.
@@ 5.6 Simulation overview
The test was run sequentially with a burn-in period and ramp (burn-in = 50 patients, ramp = 100 patients, $c_{\max}=0.6$). Under the null (no mean difference between arms), Type I error was close to or below the nominal level. Under the alternative, the implementation for continuous endpoints rejected the null with moderate power and typically crossed the threshold at an intermediate sample size (Table 5). Again, we believe that this may not be a replacement for the fixed-sample $t$-test but a conservative, anytime-valid monitoring tool that can trigger early stopping when effects are larger or clearer than anticipated.
@@ 6 Time-to-Event Outcomes
Grünwald et al. (2021) developed a safe logrank test using e-values under a proportional hazards model with a prior on the hazard ratio. We attempted to construct a nonparametric approach where validity derives solely from randomization, not from a correctly specified hazard model. We call this e-survival.
@@ 6.2 Betting Strategy
Note that unlike the binary and continuous approaches, this betting strategy uses only the sign of the cumulative score $Z_{j-1}$, not its magnitude. Once evidence favors one direction, the bet size is fixed at $\lambda_{\max}$ regardless of how strong the accumulated evidence is. This mirrors the Kelly criterion in betting theory: the optimal wager size depends on the expected edge, and $\lambda_{\max}=0.25$ is calibrated for moderate effects (HR $\approx 0.80$(Kelly, 1956). The connection between the wager, $\lambda$, and Kelly’s ideas on fraction of betting needs to be further explored. In brief, it makes sense that wager should be higher when prospects of winning are more favorable. For time-to-event, however, adapting response to events may take a long time, and a fixed wager may be preferable.